OpenStudy (anonymous):

Help on multiplying equations. I think I finally understand elimination by addition or subtraction. but How do you do it by multiplication. for instance 4x+2y=360 x+ y=105 I can't just add the two together because nothing cancels out, I have to multiply but what?

OpenStudy (anonymous):

I think I figured it out! I timesd everything in the second equation by by -2 (i kinda picked the number out of thin air) So that gave me 4x +-2y=360 -2x+-2y=-210 ---------- 2x=150 150/2= x=75 substituting 75 into one of the original equations gives me 4(75)+2y=360 300+2y=360 2y=60 y=30 so (75,30) is my soultion